Circular No. 8225

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003it IN UGC 40 Further to IAUC 8037, T. Puckett and L. Cox report the discovery of an

apparent supernova (mag 17.2) on an unfiltered CCD frame (limiting mag 19.0) taken on Oct. 13.13 UT (and confirmed on CCD frames taken by Cox on Oct. 16.07). SN 2003it, located at α = 0^{h}05^{m}48^{s}.47, δ = +27^{o}27′09′′.6 (equinox 2000.0), or 1′′.2 east and 11′′.3 north of the center of UGC 40, was not present on 1991–1993 Palomar Sky Survey images.

SUPERNOVAE 2003io, 2003iq, AND 2003is T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, write that a spectrum (range 370–750 nm) of SN 2003io (IAUC 8214), obtained by P. Berlind on Oct. 16.49 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-II supernova with well-developed hydrogen P-Cyg lines; adopting the NED recession velocity of 4010 km/s for UGC 1761, the expansion velocity derived from the minimum of the Hβ line is 7300 km/s. A spectrum of SN 2003iq (IAUC 8219), taken on Oct. 16.47, shows it to be also a type-II supernova with a very blue continuum with P-Cyg lines of hydrogen and helium; adopting the NED recession velocity of 2458 km/s for NGC 772, the expansion velocity derived from the minimum of the Hβ line is 10500 km/s. A spectrum of SN 2003is (IAUC 8224), taken on Oct. 16.18, shows it to be a type-Ic supernova, probably before maximum; the spectrum is similar to that of SN 1994I, five days before maximum (Filippenko et al. 1995, Ap.J. 450, L11).

1937 UB (HERMES) A. S. Rivkin and R. P. Binzel, Massachusetts Institute of Technology,

report that 0.8–2.5-μm spectroscopy of 1937 UB (IAUC 8223), obtained with the NASA Infrared Telescope Facility 3-m telescope on Oct. 16.43 UT, reveals an S-type asteroid spectrum; assuming an albedo of 0.24 and H = 17.5 (cf. MPEC 2003-U04), the estimated diameter is 0.9 km.

COMET C/2003 T3 (TABUR) Preliminary parabolic orbital elements from MPEC 2003-U08:

T = 2004 Apr. 24.606 TT  ω = 44.^{o}761
Ω = 348.472 }2000.0
q = 1.41801 AU  i = 49.602
2003 October 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green